# Spectral properties of symmetrized AMV operators

### v2

@@ 1. Introduction
where $\mathcal{G}_{k+1}(L^{2}(X,\mu))$ is the $(k+1)$-th Grassmannian of $L^{2}(X,\mu)$, and $\tilde{E}_{r}(f)$ is the energy functional naturally associated with $\tilde{\Delta}_{r}$ (Definition 3.4). These form a non-decreasing sequence of non-negative numbers. Our first main result states that these numbers eventually correspond to isolated eigenvalues of $-\tilde{\Delta}_{r}$ when $(X,\mathsf{d},\mu)$ is compact and uniformly locally doubling (Definition 2.1).
@@
Let $(X,\mathsf{d},\mu)$ be a compact uniformly locally doubling metric measure space. For any integer $k\geq 2$, there exists $r_{k}>0$ such that for any $r\in(0,r_{k})$, the operator $-\tilde{\Delta}_{r}$ admits $k+1$ eigenvalues
@@
After the previous $L^{2}$-convergence result, we address the question of spectral convergence, that is to say, the convergence of the associated eigenvalues and eigenfunctions. In this regard we show that, for any $k\in\mathbb{N}$, the function $r\mapsto\tilde{\lambda}_{k,r}$ is bounded in a neighborhood of 00, as proved in the course of Theorem 1. This ensures that the $k$-th lowest eigenvalue of the operator $-\tilde{\Delta}_{r}$, which we denote $\lambda_{k}(-\tilde{\Delta}_{r})$, exists for small enough $r$, and that it coincides with $\tilde{\lambda}_{k,r}$. Let $f_{k,r}$ be an $L^{2}$-normalized eigenfunction of $-\tilde{\Delta}_{r}$ associated with $\lambda_{k}(-\tilde{\Delta}_{r})$. Recall that if $\partial M=\emptyset$ (resp. $\partial M\neq\emptyset$), a Laplace (resp. Neumann) eigenvalue of $(M,g)$ is a number $\mu\geq 0$ for which there exists an associated eigenfuction $f\in\mathcal{C}^{\infty}(M)$ (resp. $\mathcal{C}_{\nu}^{\infty}(M)$) of $-\Delta_{g}$, i.e. $-\mu f=\Delta_{g}f$.
@@ Acknowledgments.
Both authors are funded by the Research Foundation – Flanders (FWO) via the Odysseus II programme no. G0DBZ23N. They thank the anonymous reviewers for their constructive suggestions, and Dmitri Pavlov for clarifying a point related to Lemma 2.1.
@@ 2. Averaging-like operators
Now we prove $\alpha_{\epsilon,N}=\mu(B_{N}(z))<\infty$. If $\alpha_{\epsilon,N}<\mu(B_{N}(z))$, then $\mu(B_{N}(z)\backslash\cup_{n}B_{\epsilon}(y_{n}))>0$, where $\{y_{n}\}_{n\in\mathbb{N}}$ is a maximizer of (8). Since the measure is inner regular, there must exist some compact $K\subset B_{N}(z)\backslash\cup_{n}B_{\epsilon}(y_{n})$ such that
@@ 2.1. Averaging operator
In the next lemma, we provide a mild condition under which $A_{r}$ defines a bounded linear operator between Lebesgue spaces.
@@
Assume that there exists $r>0$ such that
$V(\cdot,r)^{-1}\in L^{1}(X,\mu).$ ($\mathrm{A}_{r}$)
Then for any $p\in[1,+\infty]$ the linear operator $A_{r}:L^{p}(X,\mu)\to L^{p}(X,\mu)$ is well-defined and bounded with
$\|A_{r}\|_{p\to p}\leq\|V(\cdot,r)^{-1}\|_{1}^{1/p}\cdot$
Moreover, $A_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is compact.
@@
The case $p=+\infty$ is obvious, hence we assume $p<+\infty$. Let $u\in L^{p}(X,\mu)$. By Jensen’s inequality, for any $x\in X$,
@@
$\leq\left(\int_{X}\frac{1}{V(x,r)}\mathop{}\!\mathrm{d}\mu(x)\right)\|u\|_{p}^{p}.$
Since
@@
we obtain that $A_{r}$ is a Hilbert-Schmidt integral operator acting on the separable space $L^{2}(X,\mu)$ (recall Lemma 2.1); in particular, it is compact [RS80, Section IV.6]. ∎
Lemma 2.3.
Assume that $(X,\mathsf{d})$ is totally bounded. Then ($\mathrm{A}_{r}$) holds for any $r>0$.
Proof.
Remark 2.4.
Notice that if $\mu(X)=+\infty$ and $M(r)<+\infty$ then ($\mathrm{A}_{r}$) cannot hold :
This happens, for instance, on $\mathbb{R}^{n}$ endowed with the Euclidean distance and the Lebesgue measure. More generally, this property holds on any locally compact topological group endowed with a left-invariant metric compatible with the Haar measure and with infinite volume (see [Str74, Lemma 1] for more details about these spaces). Moreover, if $\mu(X)<\infty$ and $m(r)>0$, then ($\mathrm{A}_{r}$) always holds :
$\mu(X)\geq\sum_{i}V(x_{i},r).$
In the next statement, we provide an alternative way to prove the compactness of $A_{r}$ from $L^{2}(X,\mu)$ to itself. This goes through the compactness of $A_{r}$ from $L^{2}(X,\mu)$ to the space of continuous functions $\mathcal{C}(X)$ which we obtain for $X$ compact and satisfying the following condition:
@@
[…]
@@ 2.2. Adjoint
For any $x,y\in X$ and $r>0$, set
$a^{*}_{r}(x,y)\coloneqq a_{r}(y,x)=\frac{1_{B_{r}(x)}(y)}{V(y,r)}\,\cdot$
Consider $u\in L^{0}(X,\mu)$ such that $v(\cdot)\coloneqq u(\cdot)/V(\cdot,r)\in L^{1}_{\text{loc}}(X,\mu)$. For any $x\in X$ and $r>0$ such that $v$ is $\mu$-integrable on $B_{r}(x)$, set
$A^{*}_{r}u(x)\coloneqq\int_{B_{r}(x)}\frac{u(y)\mathop{}\!\mathrm{d}\mu(y)}{V(y,r)}=\int_{X}a_{r}^{*}(x,y)\mathop{}\!\mathrm{d}\mu(y).$
Notice that, just like $A_{r}u(x)$ in the previous subsection, $A^{*}_{r}u(x)$ may not make sense uniformly with respect to $x\in X$.
Let us begin with a simple observation.
@@
The next lemma gives a mild sufficient condition ensuring the well-posedeness and boundedness of the linear operator $A_{r}^{*}$ between $L^{p}$ spaces for $p>1$.
@@
Assume that there exists $r>0$ such that
$\left.\begin{array}[]{l}V(\cdot,r)^{-1}\in L^{\infty}(B_{r}(x),\mu)\qquad\forall x\in X,\\ \|A_{r}^{*}1\|_{\infty}<+\infty.\end{array}\right\}$ ($\mathrm{A}_{r}^{*}$)
Then for any $p\in[1,+\infty]$, the linear operator $A_{r}^{*}:L^{p}(X,\mu)\to L^{p}(X,\mu)$ is well-defined and bounded with
$\|A_{r}^{*}\|_{p\to p}\leq\|A_{r}^{*}1\|_{\infty}^{(p-1)/p}.$
Moreover, if ($\mathrm{A}_{r}$) additionally holds, then $A_{r}^{*}:L^{p}(X,\mu)\to L^{p}(X,\mu)$ is the adjoint of $A_{r}:L^{q}(X,\mu)\to L^{q}(X,\mu)$ for $q\in[1,+\infty]$ such that $11/p+1/q=$ 1, and the operator $A_{r}^{*}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is compact.
@@
For the proof of the first assertion, consider $u\in L^{\infty}(X,\mu)$. Then for any $x\in X$,
@@
Assume now that ($\mathrm{A}_{r}$) holds, and consider $u\in L^{p}(X,\mu)$ and $v\in L^{q}(X,\mu)$. Then
@@
where we used the Fubini–Tonelli to get the second equality, and the symmetry of the distance function through the equality $1_{B_{r}(x)}(y)=1_{B_{r}(y)}(x)$ to get the last one. As for the compactness of $A_{r}^{*}$, the result is a direct consequence of the Schauder theorem for compact operators which can be applied thanks to Lemma 2.2. ∎
@@
Assume that $(X,\mathsf{d})$ is totally bounded. Then ($\mathrm{A}_{r}^{*}$) holds for any $r>0$.
@@
As seen in the proof of Lemma 2.3, total boundedness implies $m(r)>0$ for any $r>0$ hence $V(\cdot,r)^{-1}\in L^{\infty}(B_{r}(x),\mu)$ for any $x\in X$. Total boundedness also implies that $\mu(X)<+\infty$, thus $\|A_{r}^{*}1\|_{\infty}\leq\mu(X)/m(r)<+\infty$. ∎
@@
Note that if $m(r)>0$ and $M(r)<+\infty$, then ($\mathrm{A}_{r}^{*}$) always holds with
@@
This notably happens on locally compact topological groups endowed with a left-invariant distance and their Haar measure, compare with Remark 2.4. Furthermore, observe that $A_{r}^{*}1\in L^{\infty}(X,\mu)$ can be seen as a weak variant of the comparability conditions introduced in [Ald18, MT23].
@@ 2.3. Symmetrization
Assume that there exists $r>0$ such that ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$) hold. Then $\tilde{A}_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is a compact self-adjoint operator such that
$\|\tilde{A}_{r}\|_{2\to 2}\leq\frac{\|V(\cdot,r)^{-1}\|_{1}^{1/2}+c(r)^{1/2}}{2}$
2.4. Doubling condition
Let us recall the classical doubling condition and its uniform local variant, see e.g. [HKST15] for more details.
Definition 2.1.
The space $(X,\mathsf{d},\mu)$ is called doubling if there exists $C>0$ such that for any $x\in X$ and $r>0$,
$V(x,2r)\leq CV(x,r).$ (24)
It is called uniformly uniformly locally doubling if there exist $C,r_{0}>0$ such that (24) holds for any $x\in X$ and $r\in(0,r_{0})$.
It is worth pointing out that the celebrated Bishop–Gromov theorem, see e.g. [Cha95, Theorem III.4.5], implies that any complete Riemannian manifold with a uniform lower bound on the Ricci curvature is uniformly uniformly locally doubling. This is also true for metric spaces with generalized sectional curvature bounded from below in the sense of Alexandrov [BBI01, Theorem 10.6.6] and $\mathrm{CD}(K,N)$ metric measure space [Vil09, Corollary 30.14].

### v3

@@ 1. Introduction
where $\mathcal{G}_{k+1}(L^{2}(X,\mu))$ is the $(k+1)$-th Grassmannian of $L^{2}(X,\mu)$, and $\tilde{E}_{r}(f)$ is the energy functional naturally associated with $\tilde{\Delta}_{r}$ (Definition 3.10). These form a non-decreasing sequence of non-negative numbers. Our first main result states that these numbers eventually correspond to isolated eigenvalues of $\tilde{\Delta}_{r}$ when $(X,\mathsf{d},\mu)$ is compact and uniformly locally doubling (Definition 2.8).
@@
Let $(X,\mathsf{d},\mu)$ be a compact uniformly locally doubling metric measure space. For any integer $k\geq 2$, there exists $r_{k}>0$ such that for any $r\in(0,r_{k})$, the operator $\tilde{\Delta}_{r}$ admits $k+1$ eigenvalues
@@
After the previous $L^{2}$-convergence result, we address the question of spectral convergence, that is to say, the convergence of the associated eigenvalues and eigenfunctions. In this regard we show that, for any $k\in\mathbb{N}$, the function $r\mapsto\tilde{\lambda}_{k,r}$ is bounded in a neighborhood of 0, as proved in the course of Theorem 1. This ensures that the $k$-th lowest eigenvalue of the operator $\tilde{\Delta}_{r}$, which we denote $\lambda_{k}(-\tilde{\Delta}_{r})$, exists for small enough $r$, and that it coincides with $\tilde{\lambda}_{k,r}$. Let $f_{k,r}$ be an $L^{2}$-normalized eigenfunction of $\tilde{\Delta}_{r}$ associated with $\lambda_{k}(-\tilde{\Delta}_{r})$. Recall that if $\partial M=\emptyset$ (resp. $\partial M\neq\emptyset$), a Laplace (resp. Neumann) eigenvalue of $(M,g)$ is a number $\mu\geq 0$ for which there exists an associated eigenfuction $f\in\mathcal{C}^{\infty}(M)$ (resp. $\mathcal{C}_{\nu}^{\infty}(M)$) of $\Delta_{g}$, i.e. $\mu f=\Delta_{g}f$.
@@ Acknowledgments.
Both authors are funded by the Research Foundation – Flanders (FWO) via the Odysseus II programme no. G0DBZ23N. They thank the anonymous reviewers for their numerous and constructive suggestions, and Dmitri Pavlov for clarifying a point related to Lemma 2.1.
@@ 2. Averaging-like operators
[…]
@@ 2.1. Averaging operator
Let us also set
$a^{*}_{r}(x,y)\coloneqq a_{r}(y,x)=\frac{1_{B_{r}(x)}(y)}{V(y,r)}$
for any $x,y\in X$ and $r>0$. Consider $u\in L^{0}(X,\mu)$ such that $v(\cdot)\coloneqq u(\cdot)/V(\cdot,r)\in L^{1}_{\text{loc}}(X,\mu)$. For any $x\in X$ and $r>0$ such that $v$ is $\mu$-integrable on $B_{r}(x)$, set
$A^{*}_{r}u(x)\coloneqq\int_{B_{r}(x)}\frac{u(y)\mathop{}\!\mathrm{d}\mu(y)}{V(y,r)}=\int_{X}a_{r}^{*}(x,y)u(y)\mathop{}\!\mathrm{d}\mu(y).$
Notice that, just like $A_{r}u(x)$, $A^{*}_{r}u(x)$ may not make sense uniformly with respect to $x\in X$.
For any $r>0$, we introduce the following conditions:
$\|A_{r}^{*}1\|_{\infty}<+\infty,$ ($\mathrm{I}_{r}$)
$V(\cdot,r)^{-1}\in L^{1}(X,\mu).$ ($\mathrm{II}_{r}$)
Note that ($\mathrm{II}_{r}$) implies ($\mathrm{I}_{r}$) since
$\|A_{r}^{*}1\|_{\infty}=\sup_{x\in X}|A_{r}^{*}1(x)|=\sup_{x\in X}\int_{B_{r}(x)}\frac{\mathop{}\!\mathrm{d}\mu(y)}{V(y,r)}\leq\int_{X}\frac{\mathop{}\!\mathrm{d}\mu(y)}{V(y,r)}\,\cdot$
In the next lemma, we discuss the boundedness and the compactness of the averaging operator $A_{r}$ acting on Lebesgue spaces.
@@
Assume that there exists $r>0$ such that ($\mathrm{I}_{r}$) holds. Then for any $p\in[1,+\infty]$ the linear operator $A_{r}:L^{p}(X,\mu)\to L^{p}(X,\mu)$ is well-defined and bounded with
$\|A_{r}\|_{p\to p}\leq\|A_{r}^{*}1\|_{\infty}^{1/p}\cdot$
Moreover, if ($\mathrm{II}_{r}$) holds, then $A_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is compact.
@@
The case $p=+\infty$ is obvious and holds regardless of ($\mathrm{I}_{r}$). Let us assume that $p<+\infty$. Let $u\in L^{p}(X,\mu)$. By Jensen’s inequality, for any $x\in X$,
@@
$=\int_{X}\int_{X}\frac{1}{V(x,r)}\underbrace{1_{B_{r}(x)}(y)}_{=1_{B_{r}(y)}(x)}|u(y)|^{p}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x)$
$=\int_{X}|u(y)|^{p}\underbrace{\int_{B_{r}(y)}\frac{\mathop{}\!\mathrm{d}\mu(x)}{V(x,r)}}_{=A_{r}^{*}1(y)}\mathop{}\!\mathrm{d}\mu(y)\leq\|A_{r}^{*}1\|_{\infty}\|u\|_{p}^{p}$
where we have used the Fubini–Tonelli theorem to get the second equality and ($\mathrm{I}_{r}$) for the last inequality.
Let us now assume that ($\mathrm{II}_{r}$) holds. Since
@@
we obtain that $A_{r}$ is a Hilbert-Schmidt integral operator acting on the separable space $L^{2}(X,\mu)$ (recall Lemma 2.1); in particular, $A_{r}$ is compact [RS80, Section IV.6]. ∎
In the next statement, we provide an alternative way to prove the compactness of $A_{r}$ from $L^{2}(X,\mu)$ to itself. This goes through the compactness of $A_{r}$ from $L^{2}(X,\mu)$ to the space of continuous functions $\mathcal{C}(X)$ which we obtain for compact spaces $X$ satisfying the following condition:
@@
We start by noticing that if $u\in L^{2}(X,\mu)$, then $A_{r}(u)$ is continuous. This follows from $V(\cdot,r)^{-1}$ and $\int_{B_{r}(x)}u(y)d\mu(y)$ being continuous. The former holds by assumption. To prove the latter, assume that $\|u\|_{L^{2}(X)}=1$. Then for any $x,z\in X$,
$\left|\int_{B^{r}(x)}u(y)d\mu(y)-\int_{B^{r}(z)}u(y)d\mu(y)\right|$ $\leq\left\|1_{B_{r}(x)}-1_{B_{r}(z)}\right\|_{L^{2}(X)}\left\|u\right\|_{L^{2}(X)}$ (21)
@@ 2.2. Adjoint
Let us focus now on the boundedness and the compactness of the adjoint operator $A_{r}^{*}$. We begin with a simple observation.
@@
We continue with the next lemma which covers the case $p>1$.
@@
Assume that there exists $r>0$ such that ($\mathrm{I}_{r}$) holds. Then for any $p\in[1,+\infty]$, the linear operator $A_{r}^{*}:L^{p}(X,\mu)\to L^{p}(X,\mu)$ is well-defined and bounded with
$\|A_{r}^{*}\|_{p\to p}\leq\|A_{r}^{*}1\|_{\infty}^{(p-1)/p}$
Moreover, this operator is the adjoint of $A_{r}:L^{q}(X,\mu)\to L^{q}(X,\mu)$ for $q\in[1,+\infty]$ such that $1/p+1/q=1$. Lastly, if ($\mathrm{II}_{r}$) holds, then the operator $A_{r}^{*}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is compact.
@@
For the proof of the first assertion, consider $u\in L^{\infty}(X,\mu)$. Thanks to ($\mathrm{I}_{r}$), for $\mu$-a.e. $x\in X$,
@@
Let us prove that $A_{r}$ and $A_{r}^{*}$ are adjoint of each other. Consider $u\in L^{p}(X,\mu)$ and $v\in L^{q}(X,\mu)$. Then
@@
where we used the Fubini–Tonelli to get the second equality, and the equality $1_{B_{r}(x)}(y)=1_{B_{r}(y)}(x)$ to get the last one. As for the compactness of $A_{r}^{*}$ under ($\mathrm{II}_{r}$), this result is a direct consequence of the Schauder theorem for compact operators which can be applied thanks to Lemma 2.2. ∎
2.3. Discussion on the assumptions
Let us discuss the validity of ($\mathrm{I}_{r}$) and ($\mathrm{II}_{r}$). Recall first that : ($\mathrm{II}_{r}$) $\Rightarrow$ ($\mathrm{I}_{r}$). Both properties hold on totally bounded spaces, as seen in the next lemma.
@@ 2.3. Discussion on the assumptions
Assume that $(X,\mathsf{d})$ is totally bounded. Then ($\mathrm{II}_{r}$) (and then ($\mathrm{I}_{r}$)) holds for any $r>0$.
@@
If $\mu(X)=+\infty$ and $M(r)<+\infty$ then ($\mathrm{II}_{r}$) cannot hold :
This happens, for instance, on $\mathbb{R}^{n}$ endowed with the Euclidean distance and the Lebesgue measure. More generally, this property cannot hold on a locally compact topological group endowed with a left-invariant metric compatible with the Haar measure and with infinite volume (see [Str74, Lemma 1] for more details about these spaces).
If $\mu(X)<\infty$ and $m(r)>0$, then ($\mathrm{II}_{r}$) always holds :
$\mu(X)\geq\sum_{i}V(x_{i},r).$
Let us now focus on ($\mathrm{I}_{r}$). We show below that this condition holds on so-called doubling spaces. Let us recall this classical property and its uniform local variant, see e.g. [HKST15] for more details.
Definition 2.8.
The space $(X,\mathsf{d},\mu)$ is called globally doubling if there exists $C>0$ such that for any $x\in X$ and $r>0$,
$V(x,2r)\leq CV(x,r).$ (23)
It is called uniformly locally doubling if there exist $C,r_{0}>0$ such that (23) holds for any $x\in X$ and $r\in(0,r_{0})$.
The celebrated Bishop–Gromov theorem (see e.g. [Cha95, Theorem III.4.5]) implies that any complete Riemannian manifold with a uniform lower bound on the Ricci curvature is uniformly uniformly locally doubling, and globally doubling if the uniform bound is non-negative. This is also true for metric spaces with generalized sectional curvature bounded from below in the sense of Alexandrov [BBI01, Theorem 10.6.6] and $\mathrm{CD}(K,N)$ metric measure space [Vil09, Corollary 30.14].
The next lemma relates the uniformly local doubling condition with ($\mathrm{I}_{r}$).
Lemma 2.9.
Let $(X,\mathsf{d},\mu)$ be uniformly locally doubling with parameters $C,r_{0}$. Then ($\mathrm{I}_{r}$) holds with $\|A_{r}^{*}1\|_{\infty}\leq C$ for any $r\in(0,r_{0})$.
Proof.
For any $x\in X$ and $r\in(0,r_{0})$, the triangle inequality yields that $B_{r}(x)\subset B_{2r}(y)$ for any $y\in B_{r}(x)$. Then
$A_{r}^{*}1(x)=\fint_{B_{r}(x)}\frac{V(x,r)}{V(y,r)}\mathop{}\!\mathrm{d}\mu(y)\leq\fint_{B_{r}(x)}\frac{V(y,2r)}{V(y,r)}\mathop{}\!\mathrm{d}\mu(y)\leq C.$
Remark 2.10.
Of course if $(X,\mathsf{d},\mu)$ is globally doubling with constant $C$, then ($\mathrm{I}_{r}$) holds with $\|A_{r}^{*}1\|_{\infty}\leq C$ for any $r>0$.
Remark 2.11.
The previous result notably implies that ($\mathrm{I}_{r}$) may hold in situations where ($\mathrm{II}_{r}$) does not. This happens e.g. on a non-compact Riemannian manifold $(M,g)$ with non-negative Ricci curvature endowed with its canonical Riemannian distance $\mathsf{d}$ and volume measure $\mu$. Indeed, such a space has infinite volume, and the Bishop–Gromov theorem implies that $M(r)\leq\mathbb{V}^{n}(r)$ for any $r>0$, where $\mathbb{V}^{n}(r)$ is the Lebesgue measure of an Euclidean ball of radius $r$ in $\mathbb{R}^{n}$. From Remark 2.7, we get that $M$ cannot satisfy ($\mathrm{II}_{r}$) for any $r>0$, while it does satisfies ($\mathrm{I}_{r}$) thanks to the global doubling condition.
We conclude this discussion with two final remarks. First, if $m(r)>0$ and $M(r)<+\infty$, then ($\mathrm{I}_{r}$) always holds with
@@
This happens on locally compact topological groups endowed with a left-invariant distance and their Haar measure, compare with Remark 2.7. Secondly, ($\mathrm{I}_{r}$) can easily be seen as a weak variant of the comparability conditions introduced in [Ald18, MT23].
@@ 2.4. Symmetrization
Assume that there exists $r>0$ such that ($\mathrm{I}_{r}$) holds. Then $\tilde{A}_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is a self-adjoint operator such that
$\|\tilde{A}_{r}\|_{2\to 2}\leq\|A_{r}^{*}1\|_{\infty}^{1/2}.$
Moreover, if ($\mathrm{II}_{r}$) holds, then $\tilde{A}_{r}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is compact.